OpenStudy (anonymous):

@Arnab09 Factor 2x^2 + 7x + 5. A. (2x + 1)(x + 5) B. (2x + 5)(x + 1) C. (2x + 3)(x + 2) D. The polynomial is prime

OpenStudy (anonymous):

2x^2+7x+5= 2x^2 + 2x + 5x + 5= 2x(x+1) +5(x+1)=(x+1)(2x+5)